# LVMark: Robust Watermark for latent video diffusion models

### v2

@@ Abstract
Rapid advancements in generative models have made it possible to create hyper-realistic videos. As their applicability increases, their unauthorized use has raised significant concerns, leading to the growing demand for techniques to protect the ownership of the generative model itself. While existing watermarking methods effectively embed watermarks into image-generative models, they fail to account for temporal information, resulting in poor performance when applied to video-generative models. To address this issue, we introduce a novel watermarking method called LVMark, which embeds watermarks into video diffusion models. A key component of LVMark is a selective weight modulation strategy that efficiently embeds watermark messages into the video diffusion model while preserving the quality of the generated videos. To accurately decode messages in the presence of malicious attacks, we design a watermark decoder that leverages spatio-temporal information in the 3D wavelet domain through a cross-attention module. To the best of our knowledge, our approach is the first to highlight the potential of video-generative model watermarking as a valuable tool for enhancing the effectiveness of ownership protection in video-generative models. Project website: https://kuai-lab.github.io/lvmark2024/
@@ 1 Introduction
Video generative models have significantly enhanced their ability to create high-quality videos [54, 24, 3, 6, 44]. As these advancements increase the utility of video diffusion models, it is necessary to address the issue of model ownership infringement [12, 41, 18]. Since existing video watermarking methods embed the watermark into the video rather than the generative model [23, 39], they require costly post-processing for each video and cannot verify ownership if the generative model is stolen. As a way to establish ownership of generative models, embedding watermarks directly into diffusion models has recently collected much attention [18, 12]. Inspired by watermarking methods for image diffusion models, in this paper, we propose a method to embed watermarks into video diffusion models to establish model ownership.
Figure 1: Overview of video generation and ownership identification. Top: A binary message is embedded in the latent decoder to generate a watermarked video. Bottom: The watermark decoder extracts message from modified videos (e.g., H.264 compression, frame drops, cropping) to identify the model ownership.
To the best of our knowledge, there does not exist a watermarking method for watermarks into video-generative models, but several approaches that embed watermarks into image diffusion models have been proposed so far [12, 42]. The key challenge is how to embed a watermark with minimal degradation of generated images. The common idea is to fine-tune the latent decoder of the image generator using loss functions focused on image quality and watermark accuracy; for example, Stable Signature [12] employ LPIPS [52] to assess watermarked image quality and binary cross-entropy loss for the watermark message.
In the absence of watermarking methods for video diffusion models, applying the watermarking methods from image to video diffusion models is essential. However, watermarking methods for image diffusion models do not consider video-specific features, such as temporal consistency and robustness to video-specific attacks. When image diffusion models [29, 32, 34] are applied to video diffusion models, they fine-tune image-generative models to embed messages into each video frame using a 2D watermark decoder, which results in message embedding without considering inter-frame relationships. Moreover, these methods are vulnerable to video-specific attacks, such as H.264 compression, which is the widely used compression technology that easily removes messages embedded in the videos [23, 8, 37, 43, 20]. As these methods are trained to embed a single message into the model, they require a separate training process for each watermark message. Therefore, simply applying watermarking methods developed for image diffusion models is insufficient for video diffusion models.
To address these challenges, we introduce LVMark, a novel watermarking method specifically designed for video diffusion models. The goal of LVMark is to embed random binary messages into each frame of a video, enabling the generation of videos with different watermark messages after a single training process. To embed messages effectively while preserving video quality, we selectively modulate weight parameters in the latent decoder with minimal visual impact, enabling direct embedding of multiple random messages with single training [48]. Additionally, we design a robust watermark decoder that integrates the 3D wavelet domain which has spatio-temporal information with the RGB domain through a cross-attention module, which is shown to have the ability to effectively fuse two domains [16, 15]. To balance the trade-off between visual quality and message decoding accuracy, which is common in watermarking tasks, we propose a weighted patch loss that places more emphasis on patch areas where artifacts occur. Experimental results demonstrate that our method outperforms all the reasonable baselines. Since our method does not alter the structure of the diffusion model, it prevents anonymous users from bypassing the watermarking process. With its ability to identify AI-generated videos and verify ownership, LVMark protects ownership of the creations. Our contributions in this paper can be summarized as follows:
  • We propose a novel watermarking method that embeds watermarks into video diffusion models, to the best of our knowledge, for the first time.
@@
  • Our watermark decoder is constructed to efficiently leverage the 3D wavelet domain, capturing the spatio-temporal information necessary for robust message decoding against video-specific attacks, including H.264 compression and frame drops.
  • To balance visual quality with message decoding accuracy, we introduce a weighted patch loss that effectively reduces artifacts in localized areas.
@@ 2.1 Video Diffusion Models
Diffusion models have become a powerful generative approach, refining noisy data progressively to produce high-quality outputs [36, 34, 27, 14, 46]. Latent Video Diffusion Models (LVDMs) [4] optimize spatio-temporal features within a compact latent space, reducing computational costs. OpenSora [54] and Latte [24] utilize the DiT to model long-range temporal dependencies [28]. In contrast, Stable video diffusion [3] and VideoCrafter [6] leverage U-Net to effectively capture both spatial and temporal dependencies. DynamiCrafter [44] enhances U-net based methods by modeling dynamic scenes and complex motion [50]. Both U-Net and DiT architectures play crucial roles in advancing video generation, highlighting the need for watermark methods that can be applied to both types of video diffusion models.
@@
3D discrete wavelet transform (3D DWT) is a classical method widely used to analyze temporal information of videos. 3D DWT operates along the time, height, and width of videos to decompose a low frequency subband and several high frequency subbands: $LLL$, $LLH$, $LHL$, $LHH$, $HLL$, $HLH$, $HHL$, and $HHH$ each with a size of $F/2\times H/2\times W/2$, where $L$ and $H$ denote low and high frequency, while $F$, $H$, and $W$ are frame, height, and width of video respectively. This operation analyzes the spatio-temporal information of videos. Existing video processing methods [51, 11, 49, 26, 47] utilize this property of 3D DWT to detect temporal information. Video digital watermarking methods [2, 22, 35, 17, 21] applied 3D DWT to the video for robustness of watermark. To the best of our knowledge, no research has yet explored watermarking video diffusion models using 3D DWT, underscoring the need for investigation to enhance robustness against common video modifications.
Figure 2: Training Pipeline. The training pipeline of our method is illustrated here. Top: We fine-tune the latent decoder to embed binary messages in generated videos and train the watermark decoder to retrieve messages from distorted videos. Bottom-left: We modulate layers of the latent decoder that minimally impact visual quality to embed random messages. Bottom-right: The watermark decoder combines the RGB video with low frequency subbands from a 3D wavelet transform using cross-attention to decode the binary message.
2.3 Diffusion Model Watermark
Digital watermarking plays a crucial role in protecting intellectual property, ensuring content authenticity, and enabling ownership tracking. Recently, digital watermarking technology has been applied to diffusion models. Watermarking methods for diffusion models [42, 53, 45] embed watermarks into the generated content from diffusion models without a post-processing step. However, since they mainly focus on the image-generative models, these methods perform poorly when applied to video diffusion models (Section 5.4), which highlights the need for research on watermarking in video-generative models.
3 Preliminaries
We utilize latent video diffusion models, which consist of a latent encoder $\mathcal{E}$, a latent decoder $\mathcal{D}$, and a diffusion model, which is typically based on one of two architectures: U-Net [33] or DiT [28]. The latent encoder $\mathcal{E}(v)$ maps the video $v\in\mathbb{R}^{d_{v}}$ to a latent representation $z\in\mathbb{R}^{d_{z}}$, where $d_{v}$ and $d_{z}$ denote the dimensions of video $v$ and latent vector $z$. In the forward process, noise is added to $z$, resulting in $z_{t}$, which is then denoised in the backward process to recover $z$, conditioned on inputs (e.g., text prompts). Finally, the latent decoder $\mathcal{D}$ converts $z$ back to the video $\hat{v}=\mathcal{D}(z)$. U-Net-based diffusion models focus on local features through downsampling and upsampling layers. In contrast, DiT-based models capture global dependencies and long-range correlations, making them effective for modeling complex temporal relationships. While image diffusion model watermarking methods [12, 41, 31] target U-Net-based models, we introduce a novel watermarking method applicable to both U-Net and DiT-based video diffusion models.
@@ 4 Methods
An overview of our method is shown in Figure 2. Our method fine-tunes the latent decoder $\mathcal{D}$ to generate video frames while embedding invisible binary messages into each frame.
@@ 4.1 Importance-based Weight Modulation
Our method aims to embed random messages into video diffusion models, enabling the generation of videos with different watermark messages. To achieve this, inspired by Yu et al. [48], we modulate the weight parameters of pre-trained latent decoder $\mathcal{D}$ based on importance of layers, as shown in bottom-left of Figure 2. For every iteration, we first transform the random message $m$ into the mapped message $m^{\prime}$ with the mapping network $\mathcal{M}(m):\mathbb{R}^{d_{m}}\rightarrow\mathbb{R}^{d_{m^{\prime}}}$, which is composed of two fully connected layers. Then, we modulate the $i$-th weight parameters of decoder $\mathcal{D}^{(i)}$ by performing element-wise multiplication between $\mathcal{D}^{(i)}$ and the mapped message, where the $i$-th fully connected layer $\mathcal{FC}^{(i)}(m^{\prime}):\mathbb{R}^{d_{m^{\prime}}}\rightarrow\mathbb{R}^{d_{\mathcal{D}}^{(i)}}$ is applied to match the dimension, where $d_{\mathcal{D}}^{(i)}$ is the input channel dimension of $i$-th weight parameters $\mathcal{D}^{(i)}$. Normalization is applied to ensure consistent scaling of the mapped message. The modulated weights in the $i$-th layer are defined as:
$\mathcal{D}^{(i)}_{mod}=\mathcal{D}^{(i)}\circ\big{(}\frac{\mathcal{FC}^{(i)}(\mathcal{M}(m))-\mu^{(i)}}{\sigma^{(i)}+\epsilon}\cdot\gamma^{(i)}+\beta^{(i)}\big{)},$ (1)
where $\gamma^{(i)}$ and $\beta^{(i)}$ are the $i$-th learnable parameters, while $\mu^{(i)}$ and $\sigma^{(i)}$ represent the mean and standard deviation of the message $\mathcal{FC}^{(i)}(\mathcal{M}(m))$. In our experiments, we find that modulating every layer of the decoder $\mathcal{D}$ leads to video quality degradation because it alters the parameters that have a significant impact on video generation. To minimize the undesirable impact on visual quality, we introduce importance-based weight modulation, which targets only the parameters with minimal effect on visual quality. To determine which layers to modulate, we first measure their importance. Specifically, before training, we sequentially add perturbations drawn from a normal distribution, $p\sim\mathcal{N}(0,1)$, to each layer, resulting in the perturbed weights $\mathcal{D}^{(i)}_{perturbed}=\mathcal{D}^{(i)}+p$. By evaluating the LPIPS [52] metric, which measures perceptual similarity, on videos generated with sequentially perturbed layers, our method identifies which layers have a greater impact on video quality. It modulates the least important 50% of the layers, which have less impact on the video quality.
@@ 4.2 Distortion Layer
Robustness against various video attacks is essential yet remains challenging for existing watermarking methods [23, 25]. To address this, we apply random distortions to the generated video in every iteration before extracting the message. We employ two types of distortions: spatial distortions and compression distortions. Spatial distortions refer to pixel-level distortions within each frame, such as cropping and blurring, which can be easily integrated. However, adapting to compression distortions, particularly H.264 compression, poses challenges during training due to their non-differentiable nature. To address this issue, we design an H.264 network inspired by DVMark [23] and train it to accurately replicate real H.264 compression, enhancing the robustness against H.264 compression.
4.3 Robust Video Watermark Decoder
As seen in the bottom-right of figure 2, we design a watermark decoder $\mathcal{W}$ that extracts messages by leveraging spatio-temporal information through the fusion of two domains: the generated RGB video $\hat{v}$ and its 3D wavelet transformed subband $\hat{v}_{freq}$. We use only the low frequency subband $LLL$ among the 3D wavelet subbands to enhance robustness against video compression, as various compression methods discard high-frequency features [19].
The watermark decoder is composed of two primary components: a feature extraction module and a feature fusion module. Before combining RGB domain and frequency domain, we first extract each domain’s features independently. Using six 3D convolutional blocks, we extract wavelet features $F_{freq}\in\mathbb{R}^{B\times 2048\times F\times H\times W}$ from the low frequency subband $LLL$, where $B$, $F$, $H$, and $W$ denote batch, frame, height, and width, respectively. On the other hand, we employ the ResNet50 model to extract spatial features $F_{rgb}\in\mathbb{R}^{B\times 2048\times F\times H\times W}$ from the RGB video $\hat{v}$.
$CA(F_{freq},F_{rgb})=\text{softmax}(\frac{Q\cdot K^{T}}{\sqrt{d_{k}}})\cdot V.$ (2)

### v3

@@ Abstract
Rapid advancements in video diffusion models have enabled the creation of realistic videos, raising concerns about unauthorized use and driving the demand for techniques to protect model ownership. Existing watermarking methods, while effective for image diffusion models, do not account for temporal consistency, leading to degraded video quality and reduced robustness against video distortions. To address this issue, we introduce LVMark, a novel watermarking method for video diffusion models. We propose a new watermark decoder tailored for generated videos by learning the consistency between adjacent frames. It ensures accurate message decoding, even under malicious attacks, by combining the low-frequency components of the 3D wavelet domain with the RGB features of the video. Additionally, our approach minimizes video quality degradation by embedding watermark messages in layers with minimal impact on visual appearance using an importance-based weight modulation strategy. We optimize both the watermark decoder and the latent decoder of diffusion model, effectively balancing the trade-off between visual quality and bit accuracy. Our experiments show that our method embeds invisible watermarks into video diffusion models, ensuring robust decoding accuracy with 512-bit capacity, even under video distortions.
@@ 1 Introduction
Video diffusion models have made significant progress in generating high-quality videos [45, 22, 3, 5, 37]. With these advancements, the need for methods to establish ownership of both the diffusion model and its creations has emerged. This issue can be addressed by embedding watermarks into the video diffusion model and enabling tracking.
Figure 1: Overview of video generation and ownership identification. Top: An authorized user generates a watermarked video. Bottom: Despite distortions applied to the video or model, the watermark decoder can reliably identify the video’s owner.
Existing watermarking methods applicable to video generation can be categorized into three types. The first approach uses post-processing techniques [46, 30, 23, 44]. While these methods enable accurate bit extraction, they can be easily bypassed if the diffusion model is stolen, as the watermark information is not part of the model. The second approach embeds watermarks into the latent noise of the diffusion model [35, 12]. This approach modifies the latent noise to generate altered content and requires the exact inverse process of the diffusion model for message extraction. Moreover, user-controlled embedding allows fingerprints to be excluded if the model is stolen [16]. The third approach fine-tunes a pre-trained diffusion model to embed the watermark [10, 16], ensuring watermark insertion cannot be bypassed while preserving the original content.
However, fine-tuning-based methods have only been explored for image diffusion models, not for videos. While 2D watermark decoders can be applied to individual frames, following the strategy of image diffusion watermarking methods, they fail to preserve temporal consistency and leverage the large data volume of videos for high-capacity watermarking. This underscores the need for a video-specific watermark decoder for video diffusion models.
To the best of our knowledge, this paper proposes the first fine-tuning-based watermarking method for video diffusion models, LVMark. We propose a new watermark decoder tailored for video diffusion models, focusing on temporal consistency. Extracting watermarks from the RGB domain forces the model to embed them in visible color regions, which degrades image quality and makes the watermark vulnerable to distortions. To address this, we integrate the 3D wavelet and RGB domains, embedding watermarks in the invisible frequency domain, with the cross-attention module [14, 13] enabling robust decoding. Additionally, to improve compression robustness, we use the low-frequency subband of the 3D wavelet domain, as video compression methods like H.264 tend to remove high-frequency regions. We optimize the watermark decoder alongside the latent decoder of the diffusion model by iteratively embedding and extracting random bits, ensuring resilient bit extraction despite weight variations.
An effective method to embed random watermarks in the diffusion model is weight modulation [16], which directly embeds watermarks into the model’s weight parameters. Despite the strong watermark embedding capability of this strategy, our experiments reveal that modulating all layers of the latent decoder degrades visual quality. To preserve video quality, we selectively modulate the latent decoder’s weight parameters with minimal visual impact. We evaluate the impact of each layer on video quality by injecting Gaussian noise into individual layers and measuring the LPIPS score [43]. To further enhance both visual quality and message decoding accuracy, we propose a weighted patch loss that focuses on regions where artifacts are more likely to occur, effectively balancing these two aspects.
Experimental results demonstrate that our method outperforms all reasonable baselines, particularly in terms of video consistency as well as robustness against attacks. Specifically, embedding a 512-bit watermark into DynamiCrafter [37] achieves a Fréchet Video Distance of 68.94, proving superior temporal consistency over state-of-the-art (SOTA) methods. After applying H.264 compression, crop, and frame drop attacks, our method achieves a bit accuracy of 99.16%, outperforming the SOTA method by 37.88%. We can establish model ownership even in cases of unauthorized use, as depicted in Figure 1, thanks to our method’s ability to embed high-capacity, imperceptible watermarks into video diffusion models.
Our contributions can be summarized as follows:
  • We introduce a robust watermarking method for video diffusion models.
  • We propose a watermark decoder that leverages low-frequency 3D wavelet components for robust spatio-temporal message decoding, ensuring resilience to video-specific attacks like H.264 compression and frame drops.
@@
  • To balance visual quality and message decoding accuracy, we introduce a weighted patch loss that effectively reduces localized artifacts.
@@ 2.1 Video Diffusion Models
Diffusion models have become a powerful generative approach, refining noisy data progressively to produce high-quality outputs [31, 28, 25, 11, 38]. Latent Video Diffusion Models (LVDMs) [4] optimize spatio-temporal features within a latent space, reducing computational costs. OpenSora [45] and Latte [22] utilize the DiT [26] to model long-range temporal dependencies. In contrast, Stable video diffusion [3] and VideoCrafter [5] leverage U-Net [27] to effectively capture both spatial and temporal dependencies. DynamiCrafter [37] enhances U-Net by modeling dynamic scenes and complex motion [41]. Both U-Net and DiT architectures play crucial roles in advancing video generation, highlighting the need for watermark methods that can be applied to both types of video diffusion models.
2.2 Diffusion Model Watermark
Digital watermarking plays a crucial role in protecting intellectual property, ensuring content authenticity, and enabling ownership tracking. Watermarking methods for diffusion models [36, 16, 10] embed watermarks into the generated content from diffusion models without a post-processing step. However, since they mainly focus on image diffusion models, these methods perform poorly when applied to video diffusion models (Section 4.4). Although VideoShield [12] is the first to propose watermarking in video diffusion models, it alters the generated video by embedding the watermark in latent noise and requires the inverse process of the diffusion model for message decoding. These limitations highlight the need for watermarking techniques that do not alter the generated image and allow for a simple decoding process.
Figure 2: Training pipeline. The training pipeline of our method is illustrated here. Top: We fine-tune the latent decoder to embed binary messages in generated videos and train the watermark decoder to retrieve messages from distorted videos. Bottom-left: We modulate layers of the latent decoder that minimally impact visual quality to embed random messages. Bottom-right: The watermark decoder combines the RGB video with low-frequency subbands from a 3D wavelet transform using cross-attention to decode the binary message.
@@ 2.2 Diffusion Model Watermark
3D discrete wavelet transform (3D DWT) is a classical method widely used to analyze temporal information of videos. 3D DWT operates along the time, height, and width of videos to decompose a low-frequency subband and several high-frequency subbands: $LLL$, $LLH$, $LHL$, $LHH$, $HLL$, $HLH$, $HHL$, and $HHH$ each with a size of $F/2\times H/2\times W/2$, where $L$ and $H$ denote low and high-frequency, while $F$, $H$, and $W$ are frame, height, and width of video respectively. Existing video processing methods [42, 9, 40, 24, 39] leverage the properties of 3D DWT to capture temporal information, while video digital watermarking techniques [2, 19, 29, 15, 18] utilize 3D DWT to enhance the robustness of the watermark. To the best of our knowledge, no research has yet explored watermarking video diffusion models using 3D DWT, underscoring the need for investigation to enhance robustness against common video modifications.
@@ 3 Methods
An overview of our method is shown in Figure 2. We introduce a novel watermark decoder that incorporates the video’s temporal consistency (Section 3.1). We selectively modulate the weights of the latent decoder $\mathcal{D}$ to embed imperceptible random watermarks into the model (Section 3.2). The distortion layer ensures robust watermark decoding against distortions like crop, rotation, and H.264 compression (Section 3.3). By iteratively optimizing the watermark and latent decoders with random watermark messages, our method embeds an imperceptible watermark while ensuring robust extraction (Section 3.4).
Preliminaries. We utilize latent video diffusion models, which consist of a latent encoder $\mathcal{E}$, a latent decoder $\mathcal{D}$, and a diffusion model, which is typically based on one of two architectures: U-Net [27] or DiT [26]. The latent encoder $\mathcal{E}(v)$ maps the video $v\in\mathbb{R}^{d_{v}}$ to a latent representation $z\in\mathbb{R}^{d_{z}}$, where $d_{v}$ and $d_{z}$ denote the dimensions of video $v$ and latent vector $z$. In the forward process, noise is added to $z$, resulting in $z_{t}$, which is then denoised in the backward process to recover $z$, conditioned on inputs (e.g., text prompts). Finally, the latent decoder $\mathcal{D}$ converts $z$ back to the video $\hat{v}=\mathcal{D}(z)$. U-Net-based diffusion models focus on local features through downsampling and upsampling layers. In contrast, DiT-based models capture global dependencies and long-range correlations, making them effective for modeling complex temporal relationships. While image diffusion model watermarking methods [10, 35] target U-Net-based models, we introduce a novel watermarking method applicable to both U-Net and DiT-based video diffusion models.
3.1 Robust Video Watermark Decoder
We propose a new watermark decoder $\mathcal{W}$ that extracts messages from generated videos by leveraging spatio-temporal information through the fusion of two domains: the generated RGB video $\hat{v}$ and its 3D wavelet transformed subband $\hat{v}_{freq}$, as depicted in the bottom-right of Figure 2. We use only the low-frequency subband $LLL$ among the 3D wavelet subbands to enhance robustness against video compression, as various compression methods discard high-frequency features [17].
The watermark decoder is composed of two primary components: a feature extraction module and a feature fusion module. Before combining the RGB domain and frequency domain, we first extract each domain’s features independently. Using six 3D convolutional blocks, we extract wavelet features $F_{freq}\in\mathbb{R}^{B\times 2048\times F\times H\times W}$ from the low-frequency subband $LLL$, where $B$, $F$, $H$, and $W$ denote batch, frame, height, and width, respectively. On the other hand, we employ the ResNet50 model to extract spatial features $F_{rgb}\in\mathbb{R}^{B\times 2048\times F\times H\times W}$ from the RGB video $\hat{v}$.
$CA(F_{freq},F_{rgb})=\text{softmax}\left(\frac{Q\cdot K^{T}}{\sqrt{d_{k}}}\right)\cdot V.$ (1)
@@ 3.2 Importance-based Weight Modulation
To embed random messages into the diffusion model, we propose an importance-based weight modulation module inspired by WOUAF [16], which modulates the weight parameters of the pre-trained latent decoder $\mathcal{D}$. Although WOUAF successfully embeds messages into the model, our experiments show that modulating all layers degrades visual quality by altering parameters crucial for video generation. To mitigate this, we propose a novel approach that modulates only a subset of layers with lower importance, as shown in bottom-left of Figure 2. As a pre-processing step, we quantify layer importance by sequentially perturbing each layer with noise sampled from a Gaussian distribution, $p\sim\mathcal{N}(0,0.04)$, resulting in perturbed weights: $\mathcal{D}^{(i)}_{perturbed}=\mathcal{D}^{(i)}+p$. By evaluating the LPIPS [43] metric, which measures perceptual similarity, on videos generated with sequentially perturbed layers, our method identifies which layers have a greater impact on video quality.
In the training phase, we transform the random message $m$ into the mapped message $m^{\prime}$ with the mapping network $\mathcal{M}(m):\mathbb{R}^{d_{m}}\rightarrow\mathbb{R}^{d_{m^{\prime}}}$, which is composed of two fully connected layers. Then, we modulate the $i$-th selected layer $\mathcal{D}^{(i)}$ by performing element-wise multiplication between $\mathcal{D}^{(i)}$ and the mapped message, where the $i$-th fully connected layer $\mathcal{FC}^{(i)}(m^{\prime}):\mathbb{R}^{d_{m^{\prime}}}\rightarrow\mathbb{R}^{d_{\mathcal{D}}^{(i)}}$ is applied to match the dimension, where $d_{\mathcal{D}}^{(i)}$ is the input channel dimension of $i$-th layer’s weight parameters $\mathcal{D}^{(i)}$. Normalization is applied to ensure consistent scaling of the mapped message. The modulated weights in the $i$-th layer are defined as:
$\mathcal{D}^{(i)}_{mod}=\mathcal{D}^{(i)}\circ\left(\tfrac{\mathcal{FC}^{(i)}(\mathcal{M}(m))-\mu^{(i)}}{\sigma^{(i)}+\epsilon}\cdot\gamma^{(i)}+\beta^{(i)}\right),$ (2)
where $\gamma^{(i)}$ and $\beta^{(i)}$ are the $i$-th learnable parameters, while $\mu^{(i)}$ and $\sigma^{(i)}$ represent the mean and standard deviation of the message $\mathcal{FC}^{(i)}(\mathcal{M}(m))$.
@@ 3.3 Distortion Layer
Robustness against various video attacks is essential yet remains challenging for existing watermarking methods [21, 23]. To address this, we apply random distortions to the generated video in every iteration. We employ two types of distortions: spatial distortions and compression distortions. Spatial distortions refer to pixel-level distortions within each frame, such as cropping and blurring, which can be easily integrated. However, the non-differentiable nature of H.264 compression complicates training. To overcome this, we design an H.264 network inspired by DVMark [21] to mimic real compression, improving robustness.